# Online Iterative Reinforcement Learning from Human Feedback with General Preference Model

### v2

@@ Abstract
We study Reinforcement Learning from Human Feedback (RLHF) under a general preference oracle. In particular, we do not assume that there exists a reward function and the preference signal is drawn from the Bradley-Terry model as most of the prior works do. We consider a standard mathematical formulation, the reverse-KL regularized minimax game between two LLMs for RLHF under general preference oracle. The learning objective of this formulation is to find a policy so that it is consistently preferred by the KL-regularized preference oracle over any competing LLMs. We show that this framework is strictly more general than the reward-based one, and propose sample-efficient algorithms for both the offline learning from a pre-collected preference dataset and online learning where we can query the preference oracle along the way of training. Empirical studies verify the effectiveness of the proposed framework.
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) has emerged as a pivotal technique in adapting machine learning to leverage relative feedback, especially in aligning Large Language Models (LLMs) with human values and preferences (Christiano et al.,, 2017; Ziegler et al.,, 2019). Notable examples include ChatGPT (OpenAI,, 2023), Claude (Anthropic,, 2023), and Bard (Google,, 2023). The primary goal of RLHF in the context of LLMs is to adjust the responses generated by LLMs so that they are more favorably received by human evaluators.
Inspired by the standard LLM alignment workflow (Ouyang et al.,, 2022; Bai et al., 2022b, ; Touvron et al.,, 2023), we characterize an LLM by a policy $\pi$, which takes a prompt $x\in\mathcal{X}$ and produces a response $a\in\mathcal{A}$ from the distribution $\pi(\cdot|x)$. In a typical LLM training pipeline (Ouyang et al.,, 2022; Touvron et al.,, 2023; OpenAI,, 2023), the tuning process begins with a pretrained model, which is subsequently fine-tuned using specialized and instructional data to produce an initial model $\pi_{0}$. The initial model $\pi_{0}$ is then aligned with a prompt set from some distribution $x\sim d_{0}$. The key component in RLHF is the General Preference Oracle, which is mathematically defined as follows.
@@
$\small y\sim\mathrm{Ber}\big{(}\mathbb{P}(a^{1}\succ a^{2}|x,a^{1},a^{2})\big{)},$
@@
Instead of directly optimizing against the preference oracle $\mathbb{P}$, the existing prevalent RLHF framework is reward-based (Ouyang et al.,, 2022; Touvron et al.,, 2023), which consists of three steps: (1) preference data collection, (2) reward modeling, and (3) policy optimization. Specifically, the preference dataset $\mathcal{D}$ consists of multiple tuples of the form $(x,a^{1},a^{2},y)$, whose collection process can be modeled as:
@@
where $\pi_{D}^{1}$ and $\pi_{D}^{2}$ are behavior policies and are typically set as $\pi_{0}$ (Touvron et al.,, 2023; Liu et al., 2023a, ) or some powerful closed-form LLMs (Cui et al.,, 2023). The second step is reward modeling, which is the origin of the name “reward-based”. This step can be viewed as a kind of inverse RL (Ziebart et al.,, 2008), which models some difficult-to-specify goals (preferred by the human or AI evaluators) as a scalar reward signal. Specifically, the Bradley-Terry (BT) model (Bradley and Terry,, 1952), a framework widely adopted in Ouyang et al., (2022); Bai et al., 2022a ; Touvron et al., (2023); Rafailov et al., (2023); Xiong et al., (2023), assumes that there exists a ground-truth reward function $P^{*}$ and the preference model satisfies:
@@
where $\sigma(z)=1/(1+\exp(-z))$ is the sigmoid function. Then, the reward model is taken as the Maximum Likelihood Estimation (MLE) of the BT model on the preference dataset $\mathcal{D}$ (e.g., Pacchiano et al.,, 2021; Novoseller et al.,, 2020; Ouyang et al.,, 2022; Bai et al., 2022a, ; Touvron et al.,, 2023) and is used in subsequent policy optimization steps to provide a signal for algorithms like Proximal Policy Optimization (Schulman et al.,, 2017). Despite its successes, the existence of a reward function and the BT model are strong assumptions, which may not fully capture the complicated human preferences. In particular, the BT model assumes that human preference is transitive, which means that if we prefer A to B ($\mathbb{P}(A\succ B|x,A,B)>0.5$) and we prefer B to C, then it automatically holds that we prefer A to C. This assumption, however, is contradicted by evidence of intransitivity in human decision-making (Tversky,, 1969; May,, 1954). This limitation is particularly pronounced if we consider the population-level preferences, where the ultimate preference signal is aggregated across diverse human groups (May,, 1954). This may further be evidenced that in the practical RLHF, the accuracy of the learned BT model is around $70\%$ (Bai et al., 2022a, ; Touvron et al.,, 2023; Cui et al.,, 2023), suggesting the challenges in approximating the complicated human preference by BT model. While there are some recent efforts to bypass reward modeling (Rafailov et al.,, 2023; Zhao et al.,, 2023), they are still fundamentally derived from the reward-based preference model and suffer from the aforementioned issues.
Table 1: Comparison of the test accuracy between the BT-based reward model and the preference model. The reward model and preference model are trained with the same base model and preference dataset, where the details are deferred to Section 6. We evaluate the model using the Reward-Bench (Lambert et al.,, 2024).
@@
LLaMA3-8B-it Preference 98.9 62.6 88.6 95.6
In contrast, the general preference oracle defined in Definition 1 is strictly more general than the BT model and can capture a more complicated preference pattern from the definition itself. It allows an intransitive preference model and can further capture the preference feedback from AI (Bai et al., 2022b, ), with a notable example of GPT-4 (OpenAI,, 2023), which is widely used for model evaluations in practice and may more accurately reflect real user experience (Touvron et al.,, 2023; Dong et al.,, 2023; Rafailov et al.,, 2023; Xiong et al.,, 2023). Moreover, from a practical side, the preference model construction tends to be more efficient than the reward function in terms of ranking accuracy. This is evidenced by the fact that the preference model, pairRM with 0.4B parameters (Jiang et al.,, 2023), performs comparably to a LLaMA2-13B-based reward model across a diverse set of preference targets (Cui et al.,, 2023). As a case study, we train a reward model based on the Bradley-Terry (BT) model and a preference model with the same starting checkpoint Gemma-2B-it (Team et al.,, 2024) and preference dataset [^1] We remark that the mixture of the open-source preference dataset and hyper-parameters are mainly tuned for the BT model with $>2000$ A100 hours, while the preference model adopts most of them directly. Therefore, we expect that the preference model may be even better with a more refined hyper-parameter search., with results presented in Table 1 and the training details are deferred to Section 6. As we can see, the preference model achieves much higher test accuracy in the reasoning task while maintaining comparable results in other tasks. Meanwhile, the training set we use is rather limited in the reasoning data (math and coding), so the reasoning task can be viewed as an out-of-distribution task. In this sense, the preference model may also provide a better generalization compared to the reward model. The results also extend to another case study with LLaMA3-8B-instruct, where the preference model shows promising potential in the improvement of reasoning tasks. We refer interested readers to check Zhao et al., (2023); Liu et al., 2023a for further examples with similar observations. The advantage in ranking accuracy is not only directly beneficial for the algorithms that depend on ranking information (Dong et al.,, 2023; Gulcehre et al.,, 2023), but also improves the performance of algorithms derived from the reward-based framework (i.e., Bradley-Terry model), as evidenced by the results in the study of (iterative) DPO (Xiong et al.,, 2023; Hoang Tran,, 2024).
@@
  • We make the first attempt to study the theoretical learnability of RLHF under general preference oracle with KL regularization, in both the offline setting with a pre-collected preference dataset and the online setting where we can query human feedback along the way of training, which demonstrates the potential of reward-model-free learning under general preference;
@@ 2 Problem Formulation
In this section, we formulate the RLHF with general preference learning. Suppose that there exists a preference function $P^{*}:\mathcal{X}\times\mathcal{A}\times\mathcal{A}\rightarrow\mathbb{R}$ which represents the prefererence of one action $a^{1}$ over another $a^{2}$ given a prompt $x$: $P^{*}(x,a^{1},a^{2})=\mathbb{P}(a^{1}\succ a^{2}|x,a^{1},a^{2})$.
In practical applications, we want to make the resulting LLM $\pi$ close to $\pi_{0}$ (Ziegler et al.,, 2019; Ouyang et al.,, 2022; Bai et al., 2022a, ; Rafailov et al.,, 2023). Therefore, we adopt the following KL-regularized objective:
@@
One primary reason to consider the regularized target is that the constructed preference model is only locally accurate, i.e., performs well when there is little distribution shift. For instance, if the preference model is fine-tuned on a preference dataset collected by the initial model $\pi_{0}$, it improves the in-distribution generalization, but the resulting model often performs poorly out-of-distribution (Burns et al.,, 2023). Meanwhile, even if we require human labelers to give feedback along the way, the choices of the labelers may not be representative enough or the labelers can make mistakes due to limited time, attention, or care (Geng and Liu,, 2023). Moreover, the KL divergence in the target ensures that the resulting policy is stochastic instead of deterministic (given a suitable initial checkpoint), thereby more accurately reflecting the dynamics of generative language models.
We choose $P^{*}$ as the target mostly for historical reasons (Dudík et al.,, 2015; Wang et al., 2023b, ). A choice is the relative preference $\log(P^{*}(x,a^{1},a^{2})/(1-P^{*}(x,a^{1},a^{2})))$, which is equal to $R^{*}(x,a^{1})-R^{*}(x,a^{2})$ when the BT model holds so that Equation (3) becomes two decoupled regularized-reward maximization problems in this case and automatically reduces to the setting considered in the previous work Xiong et al., (2023). While we do not handle this target directly, the analysis techniques presented in this paper readily apply to it with slight modifications.
Nash Equilibrium and Best Response. Without loss of generality, we restrict our attention to the policy class $\Pi$ consisting of the policies with the same support as $\pi_{0}$ and denote the unique Nash equilibrium (known as the Minimax Winner (Simpson,, 1969; Kramer,, 1973; Fishburn,, 1984) or the von Neumann Winner (Dudík et al.,, 2015)) as the solution of the following minimax problem as:
@@
We assume that $\mathcal{P}$ is finite and the capacity of the class is large enough so that $P^{*}\in\mathcal{P}$.
Note that the finite class assumption is just for a clear presentation and the results readily generalize to an infinite class with a bounded covering number by the standard discretization technique. We define a theoretical computation oracle as follows and defer the practical implementations to the experiment section.
@@
$\pi_{P}=\mathop{\mathrm{argmax}}_{\pi^{1}\in\Pi}\min_{\pi^{2}\in\Pi}\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a^{1}\sim\pi^{1},a^{2}\sim\pi^{2}}\Big{[}P(x,a^{1},a^{2})-\eta^{-1}D_{\mathrm{KL}}(\pi^{1}(\cdot|x)\|\pi_{0}(\cdot|x))+\eta^{-1}D_{\mathrm{KL}}(\pi^{2}(\cdot|x)\|\pi_{0}(\cdot|x))\Big{]}.$
Learning Objective. The learning objective is to find an $\epsilon$-approximate Nash policy $\hat{\pi}^{1}$ for the max-player:
@@
which means that the max-player is consistently preferred by the KL-regularized preference in the face of any competing policy $\pi^{\prime}$ up to a relaxation of $\epsilon$. To stress the non-symmetric structures of the two players, we refer to the max-player as the main agent, which aims to find her $\epsilon$-approximate Nash policy, and refer the min-player to as the enhancer, which is designed to facilitate the main agent’s learning. In particular, when $\eta$ is large enough so that the KL is roughly omitted, then, we can further obtain that
@@
In this case, the obtained policy $\hat{\pi}_{1}$ is consistently preferred by the preference oracle $P^{*}$ against any competing policies. We mention in passing that the KL penalty coefficient $\eta>0$ exhibits a trade-off between being preferred by the oracle $P^{*}$ and staying close to the initial model $\pi_{0}$, and reflects the degree of our belief in the oracle $P^{*}$. In practice, $\eta$ is typically treated as a hyper-parameter and is adjusted by parameter search (Huggingface,, 2023).
@@
$J(\dagger,\hat{\pi}^{2})-J(\hat{\pi}^{1},\dagger)=$ $J(\dagger,\hat{\pi}^{2})-J(\pi_{*}^{1},\pi_{*}^{2})+J(\pi_{*}^{1},\pi_{*}^{2})-J(\hat{\pi}^{1},\dagger)$
@@
Notations. We use the short-hand notation $\pi=(\pi^{1},\pi^{2})$ when there is no confusion. We use $P(x,\pi^{1},\pi^{2})$ to represent $\mathbb{E}_{a^{1}\sim\pi^{1},a^{2}\sim\pi^{2}}[P(x,a^{1},a^{2})]$. We use $J(x,\pi^{1},\pi^{2})$ to denote the objective function in Equation (3) without the expectation over the prompt $x\sim d_{0}$. Let $\sigma(x)$ denote the sigmoid function $1/(1+e^{-x})$. We also provide a notation table in Table 2 to improve the readability of this paper.
3 Related Work
RLHF. RLHF was first popularized in the deep RL literature by Christiano et al., (2017), which served to direct the attention of the RL community to the preference-based feedback, but may further date back to Bennett et al., (2007); Knox and Stone, (2008) in the context of machine learning. It has attracted significant attention recently, mainly due to its tremendous success in Chat-GPT (OpenAI,, 2023). The most popular and standard RLHF framework is outlined in Ouyang et al., (2022); Touvron et al., (2023) and we have described the details in Section 1. In terms of reward optimization, PPO (Schulman et al.,, 2017) is the most well-known algorithm in LLM alignment literature. However, tuning the PPO algorithm to the best performance requires extensive efforts and the result of Chat-GPT4 (OpenAI,, 2023) has not been widely reproduced so far. This motivates another line of works of algorithms that are based on supervised learning. For instance, Dong et al., (2023); Yuan et al., (2023); Touvron et al., (2023); Gulcehre et al., (2023); Ji et al., (2024) propose reward ranked finetuning, (also known as rejection sampling finetuning), which essentially learns from the best-of-n policy (Nakano et al.,, 2021) to maximize the reward. The reward-ranked finetuning algorithm is a stable policy optimization algorithm with minimal hyper-parameter configuration and was applied to the RLHF of LLaMA2 (Touvron et al.,, 2023). However, it is also observed that the reward ranked finetuning algorithm leads to considerable forgetting in a wide range of tasks (also referred to as the alignment tax), as the algorithmic design only considers reward optimization (Touvron et al.,, 2023; Lin et al.,, 2023; Chen et al.,, 2024). One approach to mitigate this issue is to use the KL-regularized formulation, which is widely adopted in the deep RL approach (e.g. PPO) (Ziegler et al.,, 2019; Wu et al.,, 2021; Ouyang et al.,, 2022; Bai et al., 2022a, ; Korbak et al.,, 2022; Li et al., 2023a, ), and other supervised-learning-based algorithms (Rafailov et al.,, 2023; Wang et al., 2023a, ; Liu et al., 2023a, ; Azar et al.,, 2023), whose theoretical property is studied in Xiong et al., (2023). Among them, (offline) Direct Preference Optimization (DPO) (Rafailov et al.,, 2023) has emerged as an attractive alternative approach to PPO with notable stability and competitive performance. Xiong et al., (2023); Hoang Tran, (2024); Yuan et al., (2024) further extend the offline DPO to the iterative (online) variant, and the resulting models demonstrate impressive performance (Hoang Tran,, 2024). However, all these algorithms are designed under the reward-based RLHF framework to maximize the underlying reward function (with appropriate regularization).
Theoretical Study of Reward-based RLHF. The theoretical study of policy optimization from preference feedback dated back to the dueling bandits (e.g., Yue et al.,, 2012; Saha,, 2021; Bengs et al.,, 2021). This was later extended to the online RL setting by Xu et al., (2020); Novoseller et al., (2020); Pacchiano et al., (2021); Chen et al., (2022), including the study of tabular online RLHF with finite state, and the study of general function approximation for capturing real-world problems with large state spaces. Zhan et al., 2023b ; Wu and Sun, (2023) further encompasses the development of reward-free learning type algorithms and posterior sampling-based algorithms tailored for online RLHF. In addition to the online setting, there is also another line of works (Zhu et al.,, 2023; Zhan et al., 2023a, ; Li et al., 2023b, ) studying the reward-based RLHF in the offline setting, which learns from a pre-determined offline dataset with suitable coverage condition over the state-action space. However, these studies only consider reward maximization and deviate from the practical applications (e.g., these frameworks admit a deterministic optimal policy). Recently, Xiong et al., (2023) first formulated the RLHF as a reverse-KL regularized contextual bandit and provided finite-sample guarantees in offline, online, and hybrid settings. We remark that all these papers consider only the reward-based RLHF framework, thus differing from ours.
Theoretical Study of RLHF under General Preference Oracle. Our work is related to Dudík et al., (2015) and Wang et al., 2023b , which also investigate preference-based RLHF under a general preference model. The major difference is that we consider the reverse-KL regularized preference, aligning closely with recent LLM advancements (Ziegler et al.,, 2019; Ouyang et al.,, 2022; Bai et al., 2022a, ; Rafailov et al.,, 2023), while Dudík et al., (2015); Wang et al., 2023b considers only the non-regularized one. Meanwhile, Dudík et al., (2015) only considers the problem of finite action, while our work and Wang et al., 2023b consider the problem with large or even infinite state-action under function approximation. In terms of learning paradigm and algorithmic design, we consider both offline learning from a pre-collected dataset and batch online learning with a sparse policy update, while Dudík et al., (2015); Wang et al., 2023b studies sequential online learning that updates policy in each step, which is not feasible in the context of LLMs. Moreover, we demonstrate that the proposed algorithms can be reasonably implemented in practice, but Dudík et al., (2015); Wang et al., 2023b only focus on information-theoretical algorithms. To summarize, compared to the previous works, the framework in this work accurately reflects real-world alignment practices thus aligning more closely with the RLHF practice.
Our work is also closely related to the IPO (Azar et al.,, 2023) and Nash learning (Munos et al.,, 2023), which also motivate new algorithmic design with a general preference oracle. We comment on the similarities and differences between our framework and theirs as follows. In terms of the problem setting, our work and Nash learning consider the minimax game under the reverse-KL regularized preference, while IPO can be interpreted to find the best response of the fixed reference policy, and may be considered as a special case of the game formulation. In terms of learning paradigm, both the IPO and Nash learning only consider learning toward a fixed and known preference oracle, and study the optimization property of the problem: how to compute the optimal policy under the given preference oracle. In contrast, we study the statistical property, where the preference model needed to be learned and our goal is to find the optimal policy under the underlying ground-truth preference model. In particular, the computational challenge is hidden in Definition 2 and Munos et al., (2023) provides a reasonable approximation of the planning oracle. In this sense, our work and Munos et al., (2023) are complementary to each other.
Finally, the concurrent work Swamy et al., (2024) studies the non-regularized general preference model in the sequential online setting and aims to find the Nash equilibrium in the context of continuous control tasks. In terms of the observation model, they assume access to the preference score $\mathbb{P}(a^{1}\succ a^{2}|x,a^{1},a^{2})$, while we only observe the preference signal $y\sim\mathrm{Ber}(\mathbb{P}(a^{1}\succ a^{2}|x,a^{1},a^{2}))$. Moreover, they design online RLHF algorithms based on a reduction to the no-regret algorithm like Hedge (Freund and Schapire,, 1997), whose techniques are fundamentally different from ours.

### v3

@@ Abstract
We investigate Reinforcement Learning from Human Feedback (RLHF) in the context of a general preference oracle. In particular, we do not assume the existence of a reward function and an oracle preference signal drawn from the Bradley-Terry model as most of the prior works do. We consider a standard mathematical formulation, the reverse-KL regularized minimax game between two LLMs for RLHF under general preference oracle. The learning objective of this formulation is to find a policy so that it is consistently preferred by the KL-regularized preference oracle over any competing LLMs. We show that this framework is strictly more general than the reward-based one, and propose sample-efficient algorithms for both the offline learning from a pre-collected preference dataset and online learning where we can query the preference oracle along the way of training. Empirical studies verify the effectiveness of the proposed framework.
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) has emerged as a pivotal technique in adapting machine learning to leverage relative feedback, especially in aligning Large Language Models (LLMs) with human values and preferences [14, 90]. Notable examples include ChatGPT [49], Claude [2], and Bard [29]. The primary goal of RLHF in the context of LLMs is to adjust the responses generated by LLMs so that they are more favorably received by human evaluators.
Inspired by the standard LLM alignment workflow [50, 5, 60], we characterize an LLM by a policy $\pi$, which takes a prompt $x\in\mathcal{X}$ and produces a response $a\in\mathcal{A}$ from the distribution $\pi(\cdot|x)$. In a typical LLM training pipeline [50, 60, 49], the tuning process begins with a pretrained model, which is subsequently fine-tuned using specialized and instructional data to produce an initial model $\pi_{0}$. The initial model $\pi_{0}$ is then aligned with a prompt set from some distribution $x\sim d_{0}$. The key component in RLHF is the General Preference Oracle, which is mathematically defined as follows.
@@
$y\sim\mathrm{Ber}\big{(}\mathbb{P}(a^{1}\succ a^{2}|x,a^{1},a^{2})$
@@
Instead of directly optimizing against the preference oracle $\mathbb{P}$, the existing prevalent RLHF framework is reward-based [50, 60], which consists of three steps: (1) preference data collection, (2) reward modeling, and (3) policy optimization. Specifically, the preference dataset $\mathcal{D}$ consists of multiple tuples of the form $(x,a^{1},a^{2},y)$, whose collection process can be modeled as:
@@
where $\pi_{D}^{1}$ and $\pi_{D}^{2}$ are behavior policies and are typically set as $\pi_{0}$ [60, 43] or some powerful closed-form LLMs [16]. The second step is reward modeling, which is the origin of the name “reward-based”. This step can be viewed as a kind of inverse RL [89], which models some difficult-to-specify goals (preferred by the human or AI evaluators) as a scalar reward signal. Specifically, the Bradley-Terry (BT) model [9], a framework widely adopted in Ouyang et al. [50], Bai et al. [4], Touvron et al. [60], Rafailov et al. [53], Xiong et al. [72], assumes that there exists a ground-truth reward function $P^{*}$ and the preference model satisfies:
@@
where $\sigma(z)=1/(1+\exp(-z))$ is the sigmoid function. Then, the reward model is taken as the Maximum Likelihood Estimation (MLE) of the BT model on the preference dataset $\mathcal{D}$ [e.g., 51, 48, 50, 4, 60] and is used in subsequent policy optimization steps to provide a signal for algorithms like Proximal Policy Optimization [56]. Despite its successes, the existence of a reward function and the BT model are strong assumptions, which may not fully capture the complicated human preferences. In particular, the BT model assumes that human preference is transitive, which means that if we prefer A to B ($\mathbb{P}(A\succ B|x,A,B)>0.5$) and we prefer B to C, then it automatically holds that we prefer A to C. This assumption, however, is contradicted by evidence of intransitivity in human decision-making [62, 45]. This limitation is particularly pronounced if we consider the population-level preferences, where the ultimate preference signal is aggregated across diverse human groups [45]. This may further be evidenced that in the practical RLHF, the accuracy of the learned BT model is around $70\%$ [4, 60, 16], suggesting the challenges in approximating the complicated human preference by BT model. While there are some recent efforts to bypass reward modeling [53, 85], they are still fundamentally derived from the reward-based preference model and suffer from the aforementioned issues.
Table 1: Comparison of the test accuracy between the BT-based reward model and the preference model. The reward model and preference model are trained with the same base model and preference dataset, where the details are deferred to Section 5. We evaluate the model on Reward-Bench [39].
@@
LLaMA3-8B-it Preference 98.9 65.2 89.5 94.8
In contrast, the general preference oracle defined in Definition 1 is strictly more general than the BT model and can capture a more complicated preference pattern from the definition itself. It allows an intransitive preference model and can further capture the preference feedback from AI [5], with a notable example of GPT-4 [49], which is widely used for model evaluations in practice and may more accurately reflect real user experience [60, 18, 53, 72]. Moreover, from a practical side, the preference model construction tends to be more efficient than the reward function in terms of ranking accuracy. This is evidenced by the fact that the preference model, pairRM with 0.4B parameters [34], performs comparably to a LLaMA2-13B-based reward model across a diverse set of preference targets [16]. As a case study, we train a reward model based on the Bradley-Terry (BT) model and a preference model with the same starting checkpoint Gemma-2B-it [59] and preference dataset [^1] We remark that the mixture of the open-source preference dataset and hyper-parameters are mainly tuned for the BT model with $>2000$ A100 hours, while the preference model adopts most of them directly. Therefore, we expect that the preference model maybe even better with a more refined hyper-parameter search., with results presented in Table 1 and the training details are deferred to Section 5. As we can see, the preference model achieves much higher test accuracy in the reasoning task while maintaining comparable results in other tasks. Meanwhile, the training set we use is rather limited in the reasoning data (math and coding), so the reasoning task can be viewed as an out-of-distribution task. In this sense, the preference model may also provide a better generalization compared to the reward model. The results also extend to another case study with LLaMA3-8B-instruct, where the preference model shows promising potential in the improvement of reasoning tasks. We refer interested readers to check Zhao et al. [85], Liu et al. [43] for further examples with similar observations. The advantage in ranking accuracy is not only directly beneficial for the algorithms that depend on ranking information [18, 30], but also improves the performance of algorithms derived from the reward-based framework (i.e., Bradley-Terry model), as evidenced by the results in the study of (iterative) DPO [72, 31].
@@
  • We make the first attempt to study the theoretical learnability of RLHF under general preference oracle with KL regularization, in the offline setting with a pre-collected preference dataset and the online setting where we can query human feedback along the way of training, which demonstrates the potential of reward-model-free learning under general preference;
@@ 2 Problem Formulation
In this section, we formulate the RLHF with general preference learning. Suppose that there exists a preference function $P^{*}:\mathcal{X}\times\mathcal{A}\times\mathcal{A}\rightarrow\mathbb{R}$ which represents the prefererence of one action $a^{1}$ over another $a^{2}$ given a prompt $x$: $P^{*}(x,a^{1},a^{2})=\mathbb{P}(a^{1}\succ a^{2}|x,a^{1},a^{2})$. In practical applications, we want to make the resulting LLM $\pi$ close to $\pi_{0}$ [90, 50, 4, 53]. Therefore, we adopt the following KL-regularized objective:
@@
One primary reason to consider the regularized target is that the constructed preference model is only locally accurate, i.e., it performs well when there is little distribution shift. For instance, if the preference model is fine-tuned on a preference dataset collected by the initial model $\pi_{0}$, it improves the in-distribution generalization, but the resulting model often performs poorly out-of-distribution [10]. Meanwhile, even if we require human labelers to give feedback along the way, the choices of the labelers may not be representative enough or the labelers can make mistakes due to limited time, attention, or care [27]. Moreover, the KL divergence in the target ensures that the resulting policy is stochastic instead of deterministic (given a suitable initial checkpoint), thereby more accurately reflecting the dynamics of generative language models.
We choose $P^{*}$ as the target mostly for historical reasons [22, 65]. A choice is the relative preference $\log(P^{*}(x,a^{1},a^{2})/(1-P^{*}(x,a^{1},a^{2})))$, which is equal to $R^{*}(x,a^{1})-R^{*}(x,a^{2})$ when the BT model holds so that (3) becomes two decoupled regularized-reward maximization problems in this case and automatically reduces to the setting considered in the previous work Xiong et al. [72]. While we do not handle this target directly, the analysis techniques presented in this paper readily apply to it with slight modifications.
Nash Equilibrium and Best Response. Without loss of generality, we restrict our attention to the policy class $\Pi$ consisting of the policies with the same support as $\pi_{0}$ and denote the unique Nash equilibrium (known as the Minimax Winner [57, 38, 24] or the von Neumann Winner [22]) as the solution of the following minimax problem as:
@@
Assume that $\mathcal{P}$ is finite and the capacity of the class is large enough so that $P^{*}\in\mathcal{P}$.
The finite class assumption is for a clear presentation and the results readily generalize to an infinite class with a bounded covering number by the standard discretization technique. We define a theoretical computation oracle as follows and defer the practical implementations to the experiment section.
@@
$\pi_{P}=\mathop{\mathrm{argmax}}_{\pi^{1}\in\Pi}\min_{\pi^{2}\in\Pi}\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a^{1}\sim\pi^{1},a^{2}\sim\pi^{2}}\Big{[}P(x,a^{1},a^{2})-\eta^{-1}\log\frac{\pi^{1}(a^{1}|x)}{\pi_{0}(a^{1}|x)}+\eta^{-1}\log\frac{\pi^{2}(a^{2}|x)}{\pi_{0}(a^{2}|x)}\Big{]}.$ (5)
Learning Objective. The goal is to find an $\epsilon$-approximate Nash policy $\hat{\pi}^{1}$ for the max-player:
@@
which means that the max-player is consistently preferred by the KL-regularized preference in the face of any competing policy $\pi^{\prime}$ up to a relaxation of $\epsilon$. To stress the non-symmetric structures of the two players, we refer to the max-player as the main agent, which aims to find her $\epsilon$-approximate Nash policy, and refer to the min-player as the enhancer, which is designed to facilitate the main agent’s learning. In particular, when $\eta$ is large enough so that the KL is roughly omitted, then, we can further obtain that
@@
In this case, the obtained policy $\hat{\pi}_{1}$ is consistently preferred by the preference oracle $P^{*}$ against any competing policies. We mention in passing that the KL penalty coefficient $\eta>0$ exhibits a trade-off between being preferred by the oracle $P^{*}$ and staying close to the initial model $\pi_{0}$, and reflects the degree of our belief in the oracle $P^{*}$. In practice, $\eta$ is typically treated as a hyper-parameter and is adjusted by parameter search [32].
@@
$\small J(\dagger,\hat{\pi}^{2})-J(\hat{\pi}^{1},\dagger)=$ $J(\dagger,\hat{\pi}^{2})-J(\pi_{*}^{1},\pi_{*}^{2})+J(\pi_{*}^{1},\pi_{*}^{2})-J(\hat{\pi}^{1},\dagger)$
@@
Notations. We use the short-hand notation $\pi=(\pi^{1},\pi^{2})$ when there is no confusion. We use $P(x,\pi^{1},\pi^{2})$ to represent $\mathbb{E}_{a^{1}\sim\pi^{1},a^{2}\sim\pi^{2}}[P(x,a^{1},a^{2})]$. We use $J(x,\pi^{1},\pi^{2})$ to denote the objective function in (3) without the expectation over the prompt $x\sim d_{0}$. Let $\sigma(x)$ denote the sigmoid function $1/(1+e^{-x})$. We also provide a notation table in Table 4 to improve the readability of this paper.
Due to space constraints, the review of the related literature is deferred to Appendix 7.